# Omni-Mol: Exploring Universal Convergent Space for Omni-Molecular Tasks

### v2

@@ Acknowledgement
Appendix A Insight for Chemical Community
We visualize the routing of tokens representing different functional groups through various experts in Omni-Mol. The results clearly demonstrate that different functional groups activate distinct experts, highlighting that molecular representations differ fundamentally from pure textual semantics. This suggests the existence of intrinsic interactions, particularly between different functional groups, rather than isolated token representations. Our visualization provides strong evidence for this phenomenon, emphasizing the structured nature of molecular representations. This insight offers valuable guidance for future research on integrating molecular representations with LLMs, paving the way for more chemically informed architectures, even more powerful AI chemist.
Figure 8: Illustration of router scores in our MoE framework for tokens representing different functional groups. The numbers correspond to specific tokens associated with 9 visualized functional groups.
@@ B.1 Comprehensive Datasets Construction
(1) USPTO (USPTO, 2020). The USPTO (United States Patent and Trademark Office) dataset is a widely used large-scale chemical reaction dataset extracted and processed from US patent texts. It encompasses a diverse range of organic reaction types, including esterification, amidation, halogenation, Suzuki coupling, Buchwald–Hartwig coupling, addition reactions, condensation reactions, and redox reactions. Following Fang et al. (2024), for the Forward Reaction Prediction task, we extract data from USPTO, and split the dataset into 124,384 training instances and 1,000 test instances. Partially following Cao et al. (2024), for the Catalyst Prediction and Solvent Prediction tasks, we similarly extract data from USPTO, splitting the training/test sets into 10,079/1,015 and 67,099/7,793, respectively.
@@
USPTO_500_MT (Lu & Zhang, 2022) is a high-quality multi-task reaction prediction dataset, derived from USPTO through manual processing (including data filtering, deduplication, etc.). This subset retains the 500 most common reaction types. Following Fang et al. (2024), for the Reagent Prediction task, we split the dataset into 124,384 training instances and 1,000 test instances.
@@
USPTO_500K (Lu & Zhang, 2022), a subset of organic chemical reaction data extracted from USPTO, is widely used in chemoinformatics for the single-step retrosynthesis task. Following Fang et al. (2024), for the Retrosynthesis task, the dataset is divided into 128,684 training instances and 1,000 test instances.
@@
USPTO-Applications (Lowe, 2017) is another commonly used subset of USPTO, primarily derived from data samples in patent applications. For the Experiment Procedure Prediction task, following Liu et al. (2024c) (along with the introduction of ORD data), we split the dataset into 80% training, 10% validation and 10% test sets.
@@
(2) ChEBI-20 (Edwards et al., 2021). ChEBI-20 is derived from the ChEBI-16 (Hastings et al., 2016) dataset, with further annotations based on PubChem, forming a comprehensive database of chemical entities in the field of biochemistry. Compared to Fang et al. (2024), ChEBI-20 provides a more extensive and detailed description of chemical compounds. Therefore, for the Molecular Captioning task, following Cao et al. (2025), we split the ChEBI-20 dataset (which contains a total of 33,010 instances) into 26,420 training instances, 3295 validation instances and 3,295 test instances.
@@
(3) QM9 (Wu et al., 2018). QM9 is a subset of the GDB-17 (Ruddigkeit et al., 2012) database, focusing on quantum chemical property prediction for small organic molecules. It provides comprehensive quantum chemical attributes for molecular compounds, including spatial geometries and electronic properties, such as HOMO/LUMO energy levels obtained via DFT calculations (Kohn & Sham, 1965). In this work, we focus on the HOMO/LUMO energy levels of molecules. For the Quantum Mechanics Property Prediction task, following Fang et al. (2024), we split the dataset into 360,113 training instances and 1,987 test instances.
@@
(4) PubChem (Kim et al., 2021). PubChem is the world’s largest open-access chemical information database, focusing on chemistry, bioinformatics, and drug discovery. It provides comprehensive support for the retrieval and analysis of molecular compound data. Partiallly following Li et al. (2024b), for the Molecular Weight Prediction, LogP Prediction, Topological Polar Surface Area Prediction, and Complexity Prediction tasks, we split the dataset into 11,979/2,000, 10,673/1,785, 11,979/2,000, and 11,979/2,000 for training and test sets, respectively. Additionally, for the Description Q&A task, also following Li et al. (2024b), we split the dataset into 56,885 training instances and 10,000 test instances.
PubChemQC (Maho, 2015) is a large-scale chemical database generated through ab initio quantum chemistry calculations, with molecular compounds sourced from PubChem. Partially following Li et al. (2024b), for the SCF Energy Prediction task, we split the dataset into 623,418 training instances and 77,993 test instances.
@@
(5) RNX Yields (Schwaller et al., 2021). The RNX Yields dataset consists of the Buchwald–Hartwig reaction (Ahneman et al., 2018) dataset and the Suzuki–Miyaura reaction (Perera et al., 2018) dataset, both collected through high-throughput experimentation (HTE). It is designed to predict reaction yields for these two reaction types. Following PRESTO, we split the dataset into 9,515 training instances and 200 test instances for Yields Regression.
@@
(6) ORD (Kearnes et al., 2021). The ORD (Open Reaction Database) is an open-source database dedicated to the standardization, storage, and sharing of organic chemistry reaction data, providing a unified data schema with structured text for organic reaction datasets. Following Liu et al. (2024c) (along with the USPTO-Applications), for Experimental Procedure Prediction task, We partition the dataset into 90% for training, 10% for validation, and 10% for testing, based on the total data volume.
@@
Based on the six datasets presented above, we construct a total of 15 tasks spanning four task types, amounting to 1.8 million data samples. To the best of our knowledge, this represents the most comprehensive dataset to date in the molecular domain. The specific partitions are illustrated in Figure 9 below.
Figure 9: Domains and Tasks. We choose 15 tasks across 4 domains, including reaction, regression, description and action.
B.2 Pre-processing

### v3

@@ 7 Conclusion
Acknowledgment
We thank the University of Maryland, College Park for providing us abundant computational resources. We sincerely thank Fanding Xu for his help in verifying the case study. His professional biochemical knowledge was invaluable to this process.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims made in abstract and introduction clearly match our theoretical and empirical results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: It is provided in the Appendix and Conclusion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The complete proof is in the Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All hyperparameters are stated in the Appendix or the main paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All code and data will be publicly released.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: They are clearly explained in the Appendix and the main paper.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Statistical significance are considered.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The runtime comparison is in the Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The work conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Broader impacts of this work are discussed in the Appendix.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: The datasets and code used in this work do not pose any risk.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: They are explicitly mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: All datasets and code will be publicly released.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix
Table of Content
  • A Experiment Environments
  • B Further Details on Datasets
  • C Further Details on Model Implementation
  • D Further Details on Training
  • E Further Details on Experimental Results
  • F More Ablation Study Results
  • G Case Study
  • H Task Definition and Prompt Templates
  • I Discussion on Generalist and Specialist
  • J Discussion on Continual Learning
Appendix A Experiment Environments
In this section, we provide a summary of our experiment environment.
Software and Driver Versions. The experiments are conducted with the following key software
  • Python 3.12.1
  • Pytorch 2.5.1
  • Transformers 4.45.2
  • CUDA 12.4
Accelerators. Training Omni-Mol costs 576 $\times$ NVIDIA A100 80G GPU hours.
@@ B.1 Comprehensive Datasets Construction
(1) USPTO [66]. The USPTO (United States Patent and Trademark Office) dataset is a widely used large-scale chemical reaction dataset extracted and processed from US patent texts. It encompasses a diverse range of organic reaction types, including esterification, amidation, halogenation, Suzuki coupling, Buchwald–Hartwig coupling, addition reactions, condensation reactions, and redox reactions. Following [21], for the Forward Reaction Prediction task, we extract data from USPTO, and split the dataset into 124,384 training instances and 1,000 test instances. Partially following [11], for the Catalyst Prediction and Solvent Prediction tasks, we similarly extract data from USPTO, splitting the training/test sets into 10,079/1,015 and 67,099/7,793, respectively.
@@
USPTO_500_MT [51] is a high-quality multi-task reaction prediction dataset, derived from USPTO through manual processing (including data filtering, deduplication, etc.). This subset retains the 500 most common reaction types. Following [21], for the Reagent Prediction task, we split the dataset into 124,384 training instances and 1,000 test instances.
@@
USPTO_500K [51], a subset of organic chemical reaction data extracted from USPTO, is widely used in chemoinformatics for the single-step retrosynthesis task. Following [21], for the Retrosynthesis task, the dataset is divided into 128,684 training instances and 1,000 test instances.
@@
USPTO-Applications [50] is another commonly used subset of USPTO, primarily derived from data samples in patent applications. For the Experiment Procedure Prediction task, following [48] (along with the introduction of ORD data), we split the dataset into 80% training, 10% validation and 10% test sets.
@@
(2) ChEBI-20 [20]. ChEBI-20 is derived from the ChEBI-16 [22] dataset, with further annotations based on PubChem, forming a comprehensive database of chemical entities in the field of biochemistry. Compared to [21], ChEBI-20 provides a more extensive and detailed description of chemical compounds. Therefore, for the Molecular Captioning task, following [10], we split the ChEBI-20 dataset (which contains a total of 33,010 instances) into 26,420 training instances, 3295 validation instances and 3,295 test instances.
@@
(3) QM9 [70]. QM9 is a subset of the GDB-17 [59] database, focusing on quantum chemical property prediction for small organic molecules. It provides comprehensive quantum chemical attributes for molecular compounds, including spatial geometries and electronic properties, such as HOMO/LUMO energy levels obtained via DFT calculations [35]. In this work, we focus on the HOMO/LUMO energy levels of molecules. For the Quantum Mechanics Property Prediction task, following [21], we split the dataset into 360,113 training instances and 1,987 test instances.
@@
(4) PubChem [32]. PubChem is the world’s largest open-access chemical information database, focusing on chemistry, bioinformatics, and drug discovery. It provides comprehensive support for the retrieval and analysis of molecular compound data. Partiallly following [41], for the Molecular Weight Prediction, LogP Prediction, and Topological Polar Surface Area Prediction tasks, we split the dataset into 11,979/2,000, 10,673/1,785, and 11,979/2,000 for training and test sets, respectively. Additionally, for the Description Q&A task, also following [41], we split the dataset into 56,885 training instances and 10,000 test instances. For the IUPAC2SELFIES task, we split the training and test sets into 54,811 and 2,764 samples, respectively. For the Text Guided Molecule Generation task, we split 11,986/2,000 for training and test sets, respectively.
@@
(5) RNX Yields [60]. The RNX Yields dataset consists of the Buchwald–Hartwig reaction [3] dataset and the Suzuki–Miyaura reaction [57] dataset, both collected through high-throughput experimentation (HTE). It is designed to predict reaction yields for these two reaction types. Following PRESTO, we split the dataset into 9,515 training instances and 200 test instances for Yields Regression.
@@
(6) ORD [31]. The ORD (Open Reaction Database) is an open-source database dedicated to the standardization, storage, and sharing of organic chemistry reaction data, providing a unified data schema with structured text for organic reaction datasets. Following [48] (along with the USPTO-Applications), for the Experimental Procedure Prediction task, We partition the dataset into 90% for training, 10% for validation, and 10% for testing, based on the total data volume.
(7) ZINC [28]. ZINC ("ZINC Is Not Commercial") is an openly accessible repository of purchasable compounds engineered for structure‐based virtual screening. It is committed to offering a no‐cost, scalable platform for in silico screens. To ensure direct compatibility with leading docking engines, each entry in ZINC undergoes a specialized preprocessing workflow and is furnished with 3D conformers, the number of rotatable bonds, and other routine molecular descriptors. For the Molecule Editing task, we split the training and test sets into 218,708 and 3,579 samples, respectively.
@@
Based on the seven datasets presented above, we construct a total of 16 tasks spanning four task types, amounting to 1.4 million data samples. To the best of our knowledge, this represents the most comprehensive dataset to date in the small molecular domain.
B.2 Preprocessing